C=200+0.5(Y-T)
T=100+0.2Y
Nx=150-0.2Y
I=300
a=400

a) What is the value of government budget deficit/surplus?

b) Assume that the government introduces an insurance of 2% to the households show this will affect consumption and determine the new equilibrium level of income.
1
Expert's answer
2020-07-01T18:17:04-0400

a)   What is the value of government budget deficit/surplus?

National Income (Y) = Consumption Expenditure (C) + Investment (I) + Government Expenditure (G) + Net exports (NX)

Y = 200+0.5 (Y-T) + 300 + 400 + 150 - 0.2Y

T=100+0.2Y

Y = 1050 + 0.5 (Y- (100+0.2Y)) - 0.2Y

Y = 1050 + 0.5 (0.8Y- 100) - 0.2Y

Y = 0.4Y- 50 + 1050 - 0.2Y

Y – 0.2Y = 1000

Y = 1000 /0.8

Y = 1,250

Budget (deficit/surplus) = G – T

Government Expenditure = 400

Tax income (T)=100+0.2Y

T=100+0.2(1250)

T = 100 + 250

T = 350

Budget deficit = 400 – 350

Budget deficit = 50


b) Introduction of 2% insurance to the households

Introduction of 2% insurance will increase the consumption expenditure and consequently raise the equilibrium level of income.

Consumption expenditure

C=200+0.5(Y-T)

C=200+0.5(1250 - 350)

C=200+0.5(900)

C=200+ 450

C= 650

Introduction of insurance of 2% increases the consumption expenditure

Insurance = 2%*650

Insurance = 13

New equilibrium level of income = Ynew = C+I+G+NX

Ynew = (650+13) + 300 + 400 + 150 - 250

Ynew = 663 + 300 + 400 + 150 - 250

Ynew = 1,263